For your senior project, you would like to build a cyclotron that will accelerate protons to 10% of the speed of light. The largest vacuum chamber you can find is 50 cm in diameter.

What magnetic field strength will you need?
Answer #1
Concepts and reason

The concepts that are to be used to solve the given problem are the centripetal force in a circular motion, force on a charged particle in a magnetic field, and equilibrium of force.

Initially, derive the expression for strength of the magnetic field by using the centripetal force in a circular motion, force on a charged particle in a magnetic field. Finally, determine the strength of the magnetic field by using the relationship between the radius and magnetic field, mass, and velocity.

Fundamentals

The centripetal force in a circular motion,

F=mv2rF = \frac{{m{v^2}}}{r}

Here, F is the force, m is the mass, r is the radius, and v is the velocity.

The expression for force on a charged particle in a magnetic field is,

F=qvBsinθF = qvB\sin \theta

Here, F is the force, q is the charge, and B is the magnetic field, and θ is the angle.

The force on the due to magnetic field is,

F=qvBF = qvB

Here, FF is the force on object, qqis the charge, vv is the speed, and BB is the magnetic field.

The force on a charged particle in a magnetic field,

F=qvBsinθF = qvB\sin \theta

The velocity of the charged particle is perpendicular to the magnetic field. Hence, the value of θ is 900.

Substitute 900 for θ inF=qvBsinθF = qvB\sin \theta .

F=qvBsin(900)=qvB\begin{array}{c}\\F = qvB\sin \left( {{{90}^0}} \right)\\\\ = qvB\\\end{array}

Use the equilibrium of force and equate the centripetal force with magnetic force and solve for the magnetic field strength B

Solve the equations F=mv2rF = \frac{{m{v^2}}}{r} and F=qvBF = qvB as follows:

mv2r=qvBB=mvqr\begin{array}{c}\\\frac{{m{v^2}}}{r} = qvB\\\\B = \frac{{mv}}{{qr}}\\\end{array}

The strength of the magnetic field,

B=mvqrB = \frac{{mv}}{{qr}}

Here, B is the strength of the magnetic field, m is the mass of the proton, v is the velocity of the proton, q is the charge of the proton, r is the radius of the vacuum chamber.

The diameter of the vacuum chamber is half of the radius. Hence, the radius is,

r=50cm2=(25cm)(1m100cm)=0.25m\begin{array}{c}\\r = \frac{{50{\rm{ cm}}}}{2}\\\\ = \left( {25{\rm{ cm}}} \right)\left( {\frac{{1\;{\rm{m}}}}{{100\;{\rm{cm}}}}} \right)\\\\ = 0.{\rm{25 m}}\\\end{array}

The proton accelerates at 10% of the speed of light, hence, the speed of the proton is,

v=3×108m/s10=3×107m/s\begin{array}{c}\\v = \frac{{3 \times {{10}^8}{\rm{ m/s}}}}{{10}}\\\\ = 3 \times {10^{7{\rm{ }}}}{\rm{m/s}}\\\end{array}

The strength of the magnetic field is,

B=mvqrB = \frac{{mv}}{{qr}}

Substitute 1.67×1027kg1.67 \times {10^{ - 27}}{\rm{ kg}} for m, 3×107m/s3 \times {10^{7{\rm{ }}}}{\rm{m/s}} for v, 1.6×1019C1.6 \times {10^{{\rm{ - 19 }}}}{\rm{C}} for q, and 0.25 m for r in B=mvqrB = \frac{{mv}}{{qr}}.

B=(1.67×1027kg)(3×107m/s)(1.6×1019C)(0.25m)=1.2525T\begin{array}{c}\\B = \frac{{\left( {1.67 \times {{10}^{ - 27}}{\rm{ kg}}} \right)\left( {3 \times {{10}^{7{\rm{ }}}}{\rm{m/s}}} \right)}}{{\left( {1.6 \times {{10}^{{\rm{ - 19 }}}}{\rm{C}}} \right)\left( {0.25{\rm{ m}}} \right)}}\\\\ = 1.2525{\rm{ T}}\\\end{array}

Ans:

The strength of the magnetic field is1.2525T1.2525{\rm{ T}}.